twice one number is 28 more than a second number. when 23 is added to the second number the result is 12 less than three times the first number, what are the numbers?
Answer by mananth(16949)

one number =x
other number =y
2x=y+28
y+23=3x-12
2x=y+28
2x-y=28................(1)
y+23=3x-12
y-3x=-35................(2)
2 x -1 y = 28 .............1
-3 x + 1 y = -35 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
2 x -1 y = 28
-3 x + 1 y = -35
Add the two equations
-1 x = -7.00
/ -1
x = 7
plug value of x in (1)
2 x + -1 y = 28
14 + -1 y = 28
-1 y = 28 -14
-1 y = 14
y = -14
7 & 14 are the numbers